OpenStudy (anonymous):

Determine the type of boundary line and shading for the graph of the inequality -5x - y (less than or equal to) -2. Dashed line with shading on the side that includes the origin. Solid line with shading on the side that does not include the origin. Dashed line with shading on the side that does not include the origin. Solid line with shading on the side that includes the origin.

jimthompson5910 (jim_thompson5910):

the inequality sign is "less than or equal to", so the line is a solid line

jimthompson5910 (jim_thompson5910):

if there's an "or equal to" part added to it, then it's solid

jimthompson5910 (jim_thompson5910):

now plug in (0,0) to see if you get a true inequality -5x - y <= -2 -5(0) - 0 <= -2 0 - 0 <= -2 0 <= -2 this is FALSE, so the origin is NOT in the shaded region

OpenStudy (anonymous):

Thanks i get it now

jimthompson5910 (jim_thompson5910):

you're welcome